OpenStudy (anonymous):

Solve for c. ab + c = 5

TheSmartOne (thesmartone):

subtract ab on both sides

OpenStudy (calculusxy):

We would like to isolate c to figure out the value. Therefore, we need to have c on one side and the solution for what c equals to on the other. What we also should know is that if there is ever an expression that says \(ab\) then it means \(a \times b\). That when multiplied together will have a single product, therefore we can simply subtract that from both sides. Notice that we don't know whether the product will be a negative or a positive, but since there isn't any specific negative sign, we can assume for now that we should subtract. Now we will have from: \[\large \rm ab + c = 5\] to \[\large \rm ab - ab + c = 5 - ab \] \(\rm ab - ab\) would equal to 0 or in other words they would cancel out. So we are left with \(\large \rm c = 5 - ab\). Now if you would like to double-check then you can substitute the variables with numbers and see whether this actually works or not.